10-2: Basic Definitions WACC Klose Outfitters Inc. believes that its optimal capital structure consists of 65% common equity and 35% debt, and its tax rate is 40%. Klose must raise additional capital to fund its upcoming expansion. The firm will have $2 million of retained earnings with a cost of rs-1590. New common stock in an amount up to $9 million would have a cost of re-1796. Furthermore, Klose can raise up to $3 million of debt at an interest rate of rd 9%, and an additional $4 million of debt at rd-- 12%. The CFO estimates that a proposed expansion would require an investment of $5.0 million. What is the WACC for the last dollar raised to complete the expansion? Round your answer to two decimal places.
Answer #1

The requirement is 5 Million.

Out of this 65% is equity. So equity capital = 0.65*5 Million = 3.25 Million

Debt = 5-3.25 = 1.75 Million

Out of the equity of 3,25 Million, 2 Million would come from the retained earnings and the remaining 1.25 million would come from the new common stock issue.

The debt of 1.75 Million would come from the initial 9% debt

WACC = 2/5 *15* + 1.25/5*17% + 1.75/5*9%*(1-0.4)

We use (1-0.4) for the debt component since tax has a tax advantage and we use (1- tax rate)

WACC = 2/5 *15* + 1.25/5*17% + 1.75/5*9%*(1-0.4)

WACC = 12.14%